OpenStudy (anonymous):

what is the simplified form of 18g^6 h^6 / 12g^9 h^4

OpenStudy (unklerhaukus):

\[\frac{18g^6 h^6 }{ 12g^9 h^4}=\frac{3\times6\times g^6 h^6 }{ 2\times6\times g^9 h^4}\rightarrow\frac{3\times\cancel{6}\times \cancel {g^6} h^{\cancel6 2}}{ 2\times \cancel{6}\times g^{\cancel 9 3} \cancel{h^4}}=\frac{3h^2}{2g^3}\]